An aqueous potassium iodate stock solution is made by dissolving 6.33 mol KIO in sufficient water for the final volume of the solution to be 4.80 L.

Calculate the molarity of the stock KIO3 solution.

A 10.0 mL aliquot is removed from the described stock solution and diluted to a total volume of 100.0 mL. Calculate the molarity of the dilute solution.


1
Expert's answer
2022-03-30T02:28:02-0400

1. Molarity of stock = Mole/Volume

= 6.33/4.8 = 1.32M

2. Using priciple of dilution = C1V1 = C2V2

1.32  × 10 = C2 × 100

C2 = 10 / 1.32

= 7.58M